# Complexity of deep computations via topology of function spaces

### v1

@@ Abstract.
We use topological methods to study complexity of deep computations and limit computations. We use topology of function spaces, specifically, the classification Rosenthal compacta, to identify new complexity classes. We use the language of model theory, specifically, the concept of independence from Shelah’s classification theory, to translate between topology and computation. We use the theory of Rosenthal compacta to characterize approximablility of deep computations, both deterministically and probabilistically.
@@ Introduction
In this paper we study limit behavior of real-valued computations as the values of certain parameters of the computation model tend towards infinity, or towards zero, or towards some other fixed value, e.g., the depth of a neural network tending to infinity, or the time interval between layers of the network tending toward zero. Recently, particular cases of this situation have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [Chen:2018], Physics-Informed Neural Networks [Raissi-Perdikaris-Karniadakis:2019], and deep equilibrium models [Bai:2019], among others). In this paper, we combine ideas of topology, measure theory, and model theory to study these limit phenomena from a unified viewpoint.
@@
Informed by model theory, to each computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model. This allows us to view computations in any given computational model as elements of a space of real-valued functions, which is called the space of types of the model. The idea of embedding models of theories into their type spaces is central in model theory. In the context of this paper, the embedding of computations into spaces of types allows us to utilize the vast theory of topology of function spaces, known as $C_{p}$-theory, to obtain results about complexity of topological limits of computations. As we shall indicate next, recent classification results for spaces of functions provide an elegant and powerful machinery to classify computations according to their levels of “tameness” or “wildness”, with the former corresponding roughly to polynomial approximability and the latter to exponential approximability. The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [Todorcevic_1999_CompactSubsetsBaire]; in logic, the classification of theories developed by Shelah [Shelah:1990]; and in statistical learning, the notion of PAC learning and VC dimension pioneered by Vapkins and Chervonenkis [Vapnik-Chervonenkis:1974, Vapnik-Chervonenkis:1971].
@@
In a previous paper [alva2024approximability], we introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [Bai:2019]). There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
@@
In [alva2024approximability], we proved a new “tame vs wild” (i.e., polynomial vs exponential) dichotomy for complexity of deep computations by invoking a classical result of Grothendieck from the 50s [Grothendieck:1952]. Under our model-theoretic Rosetta stone, polynomial approximability in the sense of computation becomes identified with the notion of continuous extendability in the sense of topology, and with the notions of stability and type definability in the sense of model theory.
In this paper, we follow a more general approach, i.e., we view deep computations as pointwise limits of continuous functions. In topology, functions that arise as the pointwise limit of a sequence of continuous functions are called functions of the first Baire class, or Baire class 1 functions, or Baire-1 for short; Baire class 1 form a step above simple continuity in the hierarchy of functions studied in real analysis (Baire class 0 functions being continuous functions). Intuitively, Baire-1 functions represent functions with “controlled” discontinuities, so they are crucial in topology and set theory.
@@
We prove a new “tame vs wild” Ramsey-theoretic dichotomy for complexity of general deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT_1978_PCompactBaire], and a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [Todorcevic_1999_CompactSubsetsBaire].
@@
Todorčević’s trichotomy regards Rosenthal compacta; these are special classes of topological spaces, defined as compact spaces that can be embedded (homeomorphically identified as a subset) within the space of Baire class 1 functions on some Polish (separable, complete metric) space, under the pointwise convergence topology. Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways, and since the late 70’s, they have played a crucial role for understanding complexity of structures of functional analysis, especially Banach spaces. Todorčević’s trichotomy has been utilized to settle longstanding problems in topological dynamics and topological entropy [glasner2022tame].
@@
Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of “Non Independence Property” (known as “NIP”) in model theory, identified by Shelah [Shelah:1971, Shelah:1990], and to the concept of Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory identified by Valiant [Valiant:1984].
@@
Going beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [argyros2008rosenthal]. Argyros, Dodos and Kanellopoulos identified seven fundamental “prototypes” of separable Rosenthal compacta, and proved that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes. They showed that if a separable Rosenthal compactum is not hereditarily separable, then it must contain an uncountable discrete subspace of the size of the continuum.
@@
We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields. However, in the interest of accessibility, we do not assume previous familiarity with high-level topology or model theory, or computing. The only technical prerequisite of the paper is undergraduate-level topology and measure theory. The necessary topological background beyond undergraduate topology is covered in section 1.
@@
In section 1, we present the basic topological and combinatorial preliminaries, and in section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed). Section 3 is devoted to the classification of deep computations, and the final section, section 4, presents the probabilistic viewpoint.
Throughout the paper, we focus on classical computation; however, by refining the model-theoretic tools, the results presented here can be extended to the realm of quantum computation and open quantum systems. This extension will be addressed in a forthcoming paper.
1. General topological preliminaries: From continuity
to Baire class 1
@@ 1. General topological preliminaries: From continuity
In this section we present the preliminaries from general topology and function space theory. We include some of the proofs for completeness, but the reader familiar with these topics may skip them.
Recall that a subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets. Note that in a metrizable space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
A Polish space is a separable and completely metrizable topological space. The most important examples are the reals $\mathbb{R}$, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), and the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology). Countable products of Polish spaces are Polish; this includes spaces like $\mathbb{R}^{\mathbb{N}}$, the space of sequences of real numbers.
In this paper, we shall often discuss subspaces, and so there is a pertinent subtlety of the definitions worth mentioning: completely metrizable space is not the same as complete metric space; for an illustrative example, the interval $(0,1)$ with the metric inherited from the reals not complete, but it is Polish since it is homeomorphic to the real line. Being Polish is a topological property while being metrically complete is not.
The following result is a cornerstone of descriptive set theory, closely tied to the work of Wacław Sierpiński and Kazimierz Kuratowski, with proofs often built upon their foundations and formalized later, notably involving Stefan Mazurkiewicz’s work on complete metric spaces.
@@
A subset $A$ of a Polish space $X$ is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set. In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence. When $Y=\mathbb{R}$, we denote this collection simply as $C_{p}(X)$. A natural question is, how do topological properties of $X$ translate into $C_{p}(X)$ and vice versa? These questions, and in general the study of these spaces, are the concern of $C_{p}$-theory, an active field of research in general topology which was pioneered by A. V. Arhangel’skiĭ and his students in the 1970’s and 1980’s [Arkhangelskii:1992]. This field has found many applications in model theory and functional analysis. For a recent survey, see [tkachuk2011cp].
A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions. If $X$ nd $Y$ are topological spaces, the Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$. As above, in the special case $Y=\mathbb{R}$ we denote $B_{1}(X,Y)$ as $B_{1}(X)$. Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)$. The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles. His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $G_{\delta}$). Note that every metrizable space is perfectly normal.
The following fact was established by Baire in his 1899 thesis. A proof can be found in Section 10 of [Todorcevic_1997_TopicsTop].
@@
Moreover, if $X$ is Polish and $f\notin B_{1}(X)$, then there exists countable $D_{0},D_{1}\subseteq X$ and reals $a<b$ such that
@@
A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact. Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) have been objects of interest for researchers in Analysis and Topological Dynamics. We begin with the following well-known result. Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}>0$ such that $|f(x)|<M_{x}$ for all $f\in A$. We include a proof for the reader’s convenience:
@@
Since $A$ is pointwise bounded, for each $x\in X$, fix $M_{x}>0$ such that $|f(x)|\leq M_{x}$ for every $f\in A$.
@@
By relative countable compactness of $A$, there is an accumulation point $g\in B_{1}(X)$ of $\{f_{n}\}_{n\in\mathbb{N}}$. It is straightforward to show that since $f$ and $g$ agree on $D_{0}\cup D_{1}$, $g$ does not have a point of continuity on the closed set $\overline{D_{0}}=\overline{D_{1}}$, which contradicts Fact 1.2.
@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
In metrizable spaces, points of closure of some subset can always be approximated by points inside the set, via a convergent sequence. For more complicated spaces, such as $C_{p}(X)$, this fails in remarkable ways. To see an example, consider the Cantor space $X=2^{\mathbb{N}}$, and for each $n\in\mathbb{N}$ define $p_{n}:X\to\{0,1\}$ by $p_{n}(x)=x(n)$ for each $x\in X$. Then $p_{n}$ is continuous for each $n$, but one can show (see Chapter 1.1 of [Todorcevic_1997_TopicsTop] for details) that the only continuous functions in the closure of $\{p_{n}\}_{n\in\mathbb{N}}$ are the functions $p_{n}$ themselves; moreover, none of the subsequences of $\{p_{n}\}_{n\in\mathbb{N}}$ converge. In some sense, this example is the worst possible scenario for convergence. The topological space obtained from this closure is well-known: it is the Stone-Čech compactification of the discrete space of natural numbers, or $\beta\mathbb{N}$ for short, and it is an important object of study in general topology.
The following theorem, established by Haskell Rosenthal in 1974, is fundamental in functional analysis, and describes a sharp division in the behavior of sequences in a Banach spaces. I
Theorem 1.4 (Rosenthal’s Dichotomy, [Rosenthal:1974]).
If $X$ is Polish and $\{f_{n}\}\subseteq C_{p}(X)$ is pointwise bounded, then either $\{f_{n}\}_{n\in\mathbb{N}}$ contains a convergent subsequence or a subsequence whose closure (in $\mathbb{R}^{X}$) is homeomorphic to $\beta\mathbb{N}$.
In other words, a pointwise bounded set of continuous functions either contains a convergent subsequence, or a subsequence whose closure is essentially the same as the example mentioned in the previous paragraphs (the “wildest” possible scenario). Note that in the preceding example, the functions are trivially pointwise bounded in $\mathbb{R}^{X}$ as the functions can only take values 0 and $1$.
The genesis of Theorem 1.4 was Rosenthal’s $\ell_{1}$ theorem, which states that the only reason why Banach space can fail to have an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences) is the presence of a bounded sequence with no weakly Cauchy subsequence. The theorem is famous for connecting diverse areas of mathematics, namely, Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, we find a similar dichotomy: Either every point of closure of the set of functions will be a Baire class 1 function, or there is a sequence inside the set that behaves in the wildest possible way. The theorem is usually not phrased as a dichotomy, but rather as an equivalence:
Theorem 1.5 (“The BFT Dichotomy”. Bourgain-Fremlin-Talagrand [BFT_1978_PCompactBaire, Theorem 4G]).
@@
    AA is relatively compact in $B_{1}(X)$, i.e., $\overline{A}\subseteq B_{1}(X)$.
@@
    For every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that _nIf_n^-1(-,a]∩⋂_nIf_n^-1[b,)=.
@@
We shall say that a set $A\subseteq\mathbb{R}^{X}$ satisfies the Independence Property, or IP for short, if it satisfies the following condition: There exists every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and $a<b$ such that for every pair of disjoint sets $E,F\subseteq\mathbb{N}$, we have
$\bigcap_{n\in E}f_{n}^{-1}(-\infty,a]\cap\bigcap_{n\in F}f_{n}^{-1}[b,\infty)\neq\emptyset.$
If $A$ satisfies the negation of this condition, we will say that $A$ satisfies NIP, or that it has the NIP.
Remark 1.7.
Note that if $X$ is compact and $A\subseteq C_{p}(X)$, then $A$ satisfies the NIP if and only if for every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and for every $a<b$ there is $I\subseteq\mathbb{N}$ such that
$\bigcap_{n\in I}f_{n}^{-1}(-\infty,a]\cap\bigcap_{n\notin I}f_{n}^{-1}[b,\infty)=\emptyset.$
To summarize, the particular case of Theorem 1.5 for $X$ compact can be stated in the following way:
@@
Let $X$ be a compact Polish space. Then, for every pointwise bounded $A\subseteq C_{p}(X)$, one and exactly one of the following two conditions must hold:

### v2

@@ Abstract.
We use topological methods to study the complexity of deep computations and limit computations. We use topology of function spaces, specifically, the classification of Rosenthal compacta, to identify new complexity classes. We use the language of model theory, specifically, the concept of independence from Shelah’s classification theory, to translate between topology and computation. We use the theory of Rosenthal compacta to characterize approximability of deep computations, both deterministically and probabilistically.
@@ Introduction
In this paper we study asymptotic behavior of computations, e.g., the depth of a neural network tending to infinity, or the time interval between layers of a time-series network tending toward zero. Recently, particular cases of this concept have attracted considerable attention in deep learning research (e.g., Neural Ordinary Differential Equations [8], Physics-Informed Neural Networks [RPK19], and deep equilibrium models [4]). The formal framework introduced here provides a unified setting to study these limit phenomena from a foundational viewpoint.
@@
Informed by model theory, to each computation in a given computation model, we associate a continuous real-valued function, called the type of the computation, that describes the logical properties of this computation with respect to the rest of the model. This allows us to view computations in any given computational model as elements of a space of real-valued functions, which is called the space of types of the model. The idea of embedding models of theories into their type spaces is central in model theory. In the context of this paper, the embedding of computations into spaces of types allows us to utilize the vast theory of topology of function spaces, known as $C_{p}$-theory, to obtain results about complexity of topological limits of computations. As we shall indicate next, recent classification results for spaces of functions provide an elegant and powerful machinery to classify computations according to their levels of “tameness” or “wildness”, with the former corresponding roughly to polynomial approximability and the latter to exponential approximability. The viewpoint of spaces of types, which we have borrowed from model theory, thus becomes a “Rosetta stone” that allows us to interconnect various classification programs: In topology, the classification of Rosenthal compacta pioneered by Todorčević [TOD99]; in logic, the classification of theories developed by Shelah [SHE90]; and in statistical learning, the notions of PAC learning and VC dimension pioneered by Vapnik and Chervonenkis [VC74, VC71].
@@
In a previous paper [ADI+24], we introduced the concept of limits of computations, which we called ultracomputations (given they arise as ultrafilter limits of standard computations) and deep computations (following usage in machine learning [4]). There is a technical difference between both designations, but in this paper, to simplify the nomenclature, we will ignore the difference and use only the term “deep computation”.
@@
In [ADI+24], we proved a new “tame vs wild” (i.e., polynomially approximable vs nonpolynimially approximable) dichotomy for the complexity of deep computations by invoking a classical result of Grothendieck from the 1950s [GRO52]. Under our model-theoretic Rosetta stone, the property of polynomial approximability of computations is identified with continuous extendibility in the sense of topology, and with the notions of stability and type definability in model theory.
Simplest among deep computations are those arising as pointwise limits of (continuous) computations proper. In topology, the first Baire class, or Baire class 1 consists of functions (also called simply “Baire-1”) arising as pointwise limits of sequences of continuous functions. Intuitively, the Baire-1 class consists of functions with “controlled” discontinuities, and lies just one level of topological complexity away from the Baire class 0 which (by definition) consists of continuous functions.
@@
We prove a new “tame vs wild” Ramsey-theoretic dichotomy for complexity of general deep computations by invoking a famous paper by Bourgain, Fremlin and Talagrand from the late 70s [BFT78], and a new trichotomy for the class of “tame” deep computations by invoking an equally celebrated result of Todorčević, from the late 90s, for functions of the first Baire class [TOD99].
@@
Todorčević’s trichotomy regards Rosenthal compacta; these are special classes of topological spaces, defined as compact spaces that can be embedded (homeomorphically identified as a subset) within the space of Baire class 1 functions on some Polish (separable, completely metrizable) space, under the topology of pointwise convergence. Rosenthal compacta exhibit “topological tameness,” meaning that they behave in relatively controlled ways; since the late 70’s, they have played a crucial role in understanding the complexity of structures of functional analysis, especially Banach spaces. Todorčević’s trichotomy has been utilized to settle longstanding problems in topological dynamics and topological entropy [GM22]. It is noteworthy that Todorčević’s proof relies on sophisticated set-theoretic forcing and infinite Ramsey theory. At the time of writing this paper, decades after his original argument, no elementary proof has been found [TOD23, HT19].
@@
Through our Rosetta stone, Rosenthal compacta in topology correspond to the important concept of “Non Independence Property” (known as “NIP”) in model theory, identified by Shelah [SHE71, SHE90], and to the concept of Probably Approximately Correct learning (known as “PAC learnability”) in statistical learning theory identified by Valiant [VAL84].
@@
Going beyond Todorčević’s trichotomy, we invoke a more recent heptachotomy for Rosenthal compacta obtained by Argyros, Dodos and Kanellopoulos [ADK08]; they identify seven fundamental “prototypes” of separable Rosenthal compacta, and show that any non-metrizable separable Rosenthal compactum must contain a “canonical” embedding of one of these prototypes. They also showed that if a separable Rosenthal compactum is not hereditarily separable, then it must contain an uncountable discrete subspace of the size of the continuum.
@@
We believe that the results presented in this paper show practitioners of computation, or topology, or descriptive set theory, or model theory, how classification invariants used in their field translate into classification invariants of other fields. In the interest of accessibility, we do not assume the reader to have previous familiarity with advanced topology, model theory, or computing. The only technical prerequisites to read this paper are undergraduate-level topology and measure theory. The necessary topological background beyond undergraduate topology is covered in section 1.
@@
In section 1, we present basic topological and combinatorial preliminaries, and in section 2, we introduce the structural/model-theoretic viewpoint (no previous exposure to model theory is needed). Section 3 is devoted to the classification of deep computations.
Throughout the paper, our results pertain to classical models of computation (particularly computations involving real-valued quantities that are known and manipulated to a finite degree of precision). The final section, Section 4, introduces a probabilistic viewpoint, the development of which we intend to pursue in future research, extending the present framework to encompass non-deterministic and quantum computations.
1. General topological preliminaries: From continuity to Baire class 1
@@
In this section we present some preliminaries from general topology and function space theory. In the interest of completeness, we some proofs that may be safely skipped by readers familiar with these topics.
Recall that a subset of a topological space is $F_{\sigma}$ if it is a countable union of closed sets, and $G_{\delta}$ if it is a countable intersection of closed sets. A space is metrizable if its topology agrees with the topology induced by some metric therein. Two such metrics inducing the same topology may induce quite different properties in the category of metric spaces. For example, the interval $(0,1)$ with the usual metric (as a subset) of the reals is not complete; however, $(0,1)$ is homeomorphic to the real line, which is complete with respect to the usual metric thereon. In a metrizable space, every open set is $F_{\sigma}$; equivalently, every closed set is $G_{\delta}$.
A Polish space is a separable and completely metrizable topological space, i.e., admitting some complete metric inducing its topology. Although other (possibly incomplete) metrics may induce the same topology, being Polish is a purely topological property. One of the most important Polish spaces is the real line $\mathbb{R}$; the others include the Cantor space $2^{\mathbb{N}}$ and the Baire space $\mathbb{N}^{\mathbb{N}}$. The class of Polish spaces is closed under countable topological products; in particular, the Cantor space $2^{\mathbb{N}}$ (the set of all infinite binary sequences, endowed with the product topology), the Baire space $\mathbb{N}^{\mathbb{N}}$ (the set of all infinite sequences of naturals, also with the product topology), and the space $\mathbb{R}^{\mathbb{N}}$ of sequences of real numbers are all Polish. Recall that the product topology on these spaces is the topology of pointwise convergence: a sequence converges in the space if and only if it converges at each coordinate index.
@@
A subset of a Polish space is itself Polish in the subspace topology if and only if it is a $G_{\delta}$ set. In particular, closed subsets and open subsets of Polish spaces are also Polish spaces.
For a proof, see [ENG89, 4.3.24].
Given two topological spaces $X$ and $Y$ we denote by $C_{p}(X,Y)$ the set of all continuous functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence. The space $C_{p}(X,\mathbb{R})$ of continuous real functions on $X$ is denoted simply $C_{p}(X)$. A natural question is, how do topological properties of $X$ translate into $C_{p}(X)$ and vice versa? This general question, and the study of these spaces in general, is the concern of $C_{p}$-theory, an active field of research in general topology which was pioneered by A. V. Arhangel’skiĭ and his students in the 1970’s and 1980’s [ARK92]. This field has found many applications in model theory and functional analysis. For a recent survey, see [TKA11].
A Baire class 1 function between topological spaces is a function that can be expressed as the pointwise limit of a sequence of continuous functions. If $X$ and $Y$ are topological spaces, the space of Baire class 1 functions $f:X\rightarrow Y$ endowed with the topology of pointwise convergence is denoted $B_{1}(X,Y)$ (as above, $B_{1}(X,\mathbb{R})$ is denoted $B_{1}(X)$). Clearly, $C_{p}(X,Y)\subseteq B_{1}(X,Y)$. The Baire hierarchy of functions was introduced by René-Louis Baire in his 1899 doctoral thesis, Sur les fonctions de variables réelles. His work moved away from the 19th-century preoccupation with “pathological” functions toward a constructive classification based on pointwise limits.
A topological space $X$ is perfectly normal if it is normal and every closed subset of $X$ is a $G_{\delta}$ (equivalently, every open subset of $X$ is a $F_{\sigma}$). Every metrizable space (hence, every Polish space) is perfectly normal.
A topological space $X$ is Baire if every countable intersection of dense open sets is dense. The Baire Category Theorem states that every Hausdorff compact or completely metrizable space (hence, every Polish space) is Baire.
The following fact was established by Baire in his 1899 thesis. A proof can be found in [TOD97, Section 10].
@@
Moreover, if $X$ is Polish and $f\notin B_{1}(X)$, then there exist countable $D_{0},D_{1}\subseteq X$ and reals $a<b$ such that
@@
A subset $L$ of a topological space $X$ is relatively compact in $X$ if the closure of $L$ in $X$ is compact. Relatively compact subsets of $B_{1}(X)$ (for $X$ Polish) are of interest analysis and topological dynamics. We begin with the following well-known result. Recall that a set $A\subseteq\mathbb{R}^{X}$ of real-valued functions is pointwise bounded if for every $x\in X$ there is $M_{x}\geq 0$ (a pointwise bound at $x$) such that $|f(x)|\leq M_{x}$ for all $f\in A$. We include a proof for the reader’s convenience:
@@ 1.1. From Rosenthal’s dichotomy to the Bourgain-Fremlin-Talagrand dichotomy to Shelah’s NIP
In metrizable spaces, points of the closure of some subset can always be approximated by points in the set proper, via a convergent sequence. For more complicated spaces, such as $C_{p}$-spaces, this fails in remarkable ways. The $n$-th coordinate map $p_{n}:2^{\mathbb{N}}\to\{0,1\}$ on the Cantor space $X=2^{\mathbb{N}}$ ($\{0,1\}^{\mathbb{N}}$) is continuous for each $n\in\mathbb{N}$, and one can show (e.g., [TOD97, Chapter 1.1]) that $\{p_{n}\}_{n\in\mathbb{N}}$ is closed, but has no convergent subsequences, in $Y=C_{p}(X)$: in a sense, such example exhibits the worst failure of sequential convergence possible. The closure of $\{p_{n}\}$ in $\{0,1\}^{X}$ (or in $\mathbb{R}^{X}$ for that matter) realizes the Stone-Čech compactification $\beta\mathbb{N}$ of the discrete space of natural numbers; it is an important object of study in general topology.
The following theorem, proved by Haskell Rosenthal in 1974, is fundamental in functional analysis and captures a sharp division in the behavior of sequences in a Banach spaces.
Theorem 1.4 (Rosenthal’s Dichotomy, [ROS74]).
If $X$ is Polish and $\{f_{n}\}\subseteq C_{p}(X)$ is pointwise bounded, then $\{f_{n}\}_{n\in\mathbb{N}}$ has a convergent subsequence, or a subsequence whose closure in $\mathbb{R}^{X}$ is homeomorphic to $\beta\mathbb{N}$.
Rosenthal’s Dichotomy states that a pointwise bounded set of continuous functions contains either a convergent subsequence, or a subsequence whose closure is essentially the same as the example mentioned in the previous paragraphs (i.e., “wildest” possible). The genesis of this theorem was Rosenthal’s $\ell_{1}$-Theorem”, which states that a Banach space includes an isomorphic copy of $\ell_{1}$ (the space of absolutely summable sequences), or else every bounded sequence therein is weakly Cauchy. The $\ell_{1}$-Theorem connects diverse areas: Banach space geometry, Ramsey theory, set theory, and topology of function spaces.
As we move from $C_{p}(X)$ to the larger space $B_{1}(X)$, a dichotomy paralleling the $\ell_{1}$-Theorem holds: Either every point of the closure of a set of functions is a Baire class 1 function, or there is a sequence in the set behaving in the wildest possible way. This result is usually not phrased as a dichotomy, but rather as an equivalence as in Theorem 1.5 below.
First, we introduce some useful notation. For any set $A\subseteq\mathbb{R}^{X}$ and any real $a$, define
[…]
(In case $A=\emptyset$, we define $X^{\emptyset}_{\geq a}=X=X^{\emptyset}_{\leq a}$.) For any sequence $\{f_{n}\}\subseteq\mathbb{R}^{X}$ and $I\subseteq\mathbf{N}$, define $I^{\complement}\coloneqq\mathbb{N}\setminus I$ and $f_{I}\coloneqq\{f_{i}:i\in I\}$.
Theorem 1.5 (“The BFT Dichotomy”. Bourgain-Fremlin-Talagrand [BFT78]).
@@
    AA is relatively compact in $B_{1}(X)$.
@@
    For every $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and every $a<b$ there is $I\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{I}}\cap X_{\geq b}^{f_{I^{\complement}}}=\emptyset$.
[…]
The sets $X_{\leq a}^{f_{I}}$ and $X_{\geq b}^{f_{I^{\complement}}}$ appearing in condition Theorem 1.5(ii) are defined, respectively, in terms of $|I|$-many inequalities of the form $f_{i}(x)\leq a$, and $|I^{\complement}|$-many of the form $f_{j}(x)\geq b$. Thus, at least one of $X_{\leq a}^{f_{I}}$ and $X_{\geq b}^{f_{I^{\complement}}}$ is defined by the satisfaction of infinitely (countably) many inequalities. For our purposes, it is more natural to consider only finitely many inequalities at a time, which motivates the definitions below.
@@
We say that a function collection $A\subseteq\mathbb{R}^{X}$ has the finitary No-Independence Property (NIP) if, for all sequences $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$, there exist finite disjoint sets $E,F\subseteq\mathbb{N}$ such that $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}=\emptyset$. We say that such $E,F$ witness finitary NIP for $A$, $\{f_{n}\}$ and $a,b$.
A set $A\subseteq\mathbb{R}^{X}$ has the finitary Independence Property (IP) if it does not have finitary NIP, i.e., if there exists a sequence $\{f_{n}\}_{n\in\mathbb{N}}\subseteq A$ and reals $a<b$ such that for every pair of finite disjoint sets $E,F\subseteq\mathbb{N}$, we have $X_{\leq a}^{f_{E}}\cap X_{\geq b}^{f_{F}}\neq\emptyset$.
If the word “finite” is omitted in the above definitions, we obtain the definitions of countable NIP (weaker than finitary NIP) and countable IP (stronger than finitary IP), respectively.
If we insist on witnesses $E,F\subseteq\mathbb{N}$ such that $F=E^{\complement}$, we call the respective properties “BFT-NIP” (even weaker than countable NIP) and “BFT-IP” (even stronger than countable IP). Thus, Theorem 1.5 becomes that statement, for pointwise bounded function collections $A\subseteq C_{p}(X)$, that $A$ is relatively compact in $B_{1}(X)$ if and only if $A$ has BFT-NIP.
Unless otherwise unspecified, IP/NIP shall mean finitary IP/NIP henceforth.
Proposition 1.7.
If $X$ is compact and $A\subseteq C_{p}(X)$, then $A$ has BFT-NIP if and only if it has finitary NIP.
(No pointwise boundedness is assumed of $A$.)
Proof.
Trivially (as per the preceding discussion), finitary NIP implies BFT-NIP. Reciprocally, assume that $X$ is compact and has finitary IP. Fix $A\subseteq C_{p}(X)$, a sequence $\{f_{n}\}\subseteq A$ and reals $r<s$. For any $I,J\subseteq\mathbb{N}$ (eventually disjoint in applications), write $X_{I,J}$ for $X^{f_{I}}_{\leq r}\cap X^{f_{J}}_{\geq s}$. For $I\subseteq I^{\prime}\subseteq\mathbb{N}$ and $J\subseteq J^{\prime}\subseteq\mathbb{N}$, we have $X_{I,J}\supseteq X_{I^{\prime},J^{\prime}}$; moreover, $X_{I,J}=\bigcap_{E\subseteq I,F\subseteq J}X_{E,F}$, where the index variables $E\subseteq I$, $F\subseteq J$ range over finite subsets of $I,J$, respectively. Clearly, $E,F\subseteq\mathbb{N}$ witness finitary NIP for $\{f_{n}\}$ if and only if $X^{E,F}=\emptyset$.
Fix $I\subseteq\mathbb{N}$. Since $\{f_{n}\}\subseteq A\subseteq C_{p}(X)$ is a sequence of continuous functions, and $X$ is compact, the nested family $\{X_{E,F}:E\subseteq I,F\subseteq I^{\complement}\}$ consists of closed, thus compact, sets. Since $A$ has finitary IP by hypothesis, the nested family consists of nonempty sets, hence its intersection $X_{I,I^{\complement}}\neq\emptyset$ by compactness. This holds for arbitrary $\{f_{n}\}\subseteq A$ and $r<s$, so $A$ has BFT-IP. ∎
@@
Let $X$ be a metrizable compact (hence Polish) space. For every pointwise bounded $A\subseteq C_{p}(X)$, the following properties are all equivalent: